OpenStudy (anonymous):

2x-y=5 9x-4y=23 use method of substitution

OpenStudy (alexwee123):

solve out for either x or y in one of the equations then sub it in for the other equation

OpenStudy (anonymous):

first choose which equation you want to work with 2x-y=5 < ----- 9x-4y=23 < ----

OpenStudy (anonymous):

2x-y=5

OpenStudy (anonymous):

ok next step what variable to want to solve the equation for x or y

OpenStudy (anonymous):

y

OpenStudy (anonymous):

ok next if you you solve the equation for y what will it look like?

OpenStudy (anonymous):

5=2x-y??

OpenStudy (anonymous):

if you are solving for y y is supposed to be isolated = by itself

OpenStudy (anonymous):

ok. how would that be written?

OpenStudy (anonymous):

to solve equations you know that any operations performed one side of the equals sign must be performed on the other side right?

OpenStudy (anonymous):

so how do i get rid of this | v 2x-y=5